OpenStudy (anonymous):

Penny needs to double the size of her science project globe from a radius of 11 inches to 22 inches. How much material did Penny originally use to cover the globe? How much more material will she need to cover the globe at a radius of 22 inches?

OpenStudy (ajprincess):

the shape of a globe is sphere. to find the amount of material used find the volume of the sphere. the formula to find the volume of a sphere is V=(4/3)*pi*r^3 here r-radius. in the original case r=11 in the second case r=22. the difference between the volumes will give u the additional material needed. Does that help? @DaWizjr

OpenStudy (anonymous):

Thats the answer?

OpenStudy (anonymous):

No I need the answer so I can do the other problems

OpenStudy (ajprincess):

original case: V_o=(4/3)*pi*(11)^3 =?

OpenStudy (anonymous):

(5324 pi)/3

OpenStudy (ajprincess):

(5324 pi)/3=?

OpenStudy (anonymous):

Thats all I could get

OpenStudy (anonymous):

How much material did Penny originally use to "cover" the globe? correct me if I'm wrong, but does this imply surface area?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then your initial material would be given by the equation S= 4pir^2 I believe where r=11 then you would replace r with 22 for part b and find the difference for part c

OpenStudy (anonymous):

I just need the answer so I could do the rest

OpenStudy (anonymous):

Okay, if it's surface area then my version comes out to 4*pi*11^2 = 484 pi (part a) part b = 4*pi*22^2 = 1936pi part c = 1936pi-484 pi = 1452 pi If it means volume, then the answers would be part a= (5324 pi)/3 (as ajprincess pointed out) part b =42952pi/3 part c = 42952pi/3-(5324 pi)/3=37268pi /3

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

@ajprincess @AnElephant